The decomposition of N2O5 from the equation below

2N2O5 ----> 4NO2 + O2 has a rate of law of 1×10–³s-¹.

What is the rate of reaction when the concentration of

N2O5 is 0.0001mol/L.


1
Expert's answer
2021-05-31T05:58:09-0400

rate of reaction = k [N2O5]

rate of reaction = 1×10-3 × [0.0001]

rate of reaction = 1× 10-3 × [ 1× 10-3 ]

rate of reaction = 1 × 10-6 "molL^{-1}s^{-1}"